# Connectivity of partial basis complexes of freely decomposable groups

### v1

@@ 3. Proof of the main theorems
is at least $(n-2)$-connected. By Quillen’s fibre theorem, the map $i$ is an $(n-1)$-equivalence.
Finally, the codomain $\mathcal{FF}$ is at least $(n-2)$-connected by Theorem 1.2. Therefore $\mathcal{PB}$ is $(n-2)$-connected.
To conclude with the proof, we show the Cohen–Macaulayness. As noted in the paragraph before Theorem 1.3, we can regard $\mathcal{PB}$ as a simplicial complex. The link of a simplex $\mathcal{A}^{\prime}\in\mathcal{PB}\cup\{\emptyset\}$ can be identified with $\mathcal{PB}(A,\mathcal{A}^{\prime})$, which we saw is $(\operatorname{crk}(\mathcal{A}^{\prime})-2)$-connected. But this link has dimension $\operatorname{crk}(\mathcal{A}^{\prime})-1$, so it is spherical of this dimension. Thus $\mathcal{PB}$ is Cohen–Macaulay of dimension $n-1$. ∎
Proof of Theorem 1.1.
Let $A$ be a free group of finite rank $n$. Since $\mathcal{B}(A)$ is the inflation of $\mathcal{PB}(A,\emptyset)$, and the latter is Cohen-Macaulay of dimension $n-1$ by Theorem 1.3, it follows that $\mathcal{B}(A)$ is also Cohen–Macaulay of dimension $n-1$ (cf. Corollary 6.4 of [BWW:PosetFiberTheorems]). ∎
[…]

### v2

@@ 3. Proof of the main theorems
is at least (n2)(n-2)-connected. By Quillen’s fibre theorem, the map $i$ is an (n1)(n-1)-equivalence.
Finally, the codomain $\mathcal{FF}$ is at least (n2)(n-2)-connected by Theorem˜1.2. Therefore $\mathcal{PB}$ is (n2)(n-2)-connected.
To conclude with the proof, we show the Cohen–Macaulayness. As noted in the paragraph before Theorem˜1.3, we can regard $\mathcal{PB}$ as a simplicial complex. The link of a simplex $\mathcal{A}^{\prime}\in\mathcal{PB}\cup\{\emptyset\}$ can be identified with $\mathcal{PB}(A,\mathcal{A}^{\prime})$, which we saw is $(\operatorname{crk}(\mathcal{A}^{\prime})-2)$-connected. But this link has dimension $\operatorname{crk}(\mathcal{A}^{\prime})-1$, so it is spherical of this dimension. Thus $\mathcal{PB}$ is Cohen–Macaulay of dimension n1n-1. ∎
Before we conclude with the proof of Theorem˜1.1, we recall the notion of inflation of a simplicial complex. The following definition is a generalisation of the one given in [2], in the context of inflation of finite simplicial complexes. The adaptation presented here corresponds to Definition 4.3 in [20].
Definition 3.5.
Let $K$ be any simplicial complex (not necessarily finite) and $\mathcal{P}=(P_{v})_{v\in V(K)}$ a family of non-empty sets indexed on the vertex set $V(K)$ of $K$. The inflation of $K$ by $\mathcal{P}$, which we denote by $(K,\mathcal{P})$, is the simplicial complex whose simplices are of the form $\{(v_{0},p_{0}),\ldots,(v_{k},p_{k})\}$ such that $\{v_{0},\ldots,v_{k}\}$ is a $k$-simplex of $K$ and $p_{i}\in P_{v_{i}}$ for all $i$.
There is a natural map from the inflation to the original complex $p:(K,\mathcal{P})\to K$ that takes a simplex $\{(v_{0},p_{0}),\ldots,(v_{k},p_{k})\}\in(K,\mathcal{P})$ to $\{v_{0},\ldots,v_{k}\}\in K$. Note that $(K,\mathcal{P})$ and $K$ have the same dimension. By studying the fibres of this map, one can conclude that, if $K$ is finite-dimensional, then $(K,\mathcal{P})$ is Cohen–Macaulay if and only if $K$ is. See Proposition 5.10 of [20] for more details.
Proof of Theorem˜1.1.
Let $A$ be a free group of finite rank $n$. The complex $\mathcal{B}(A)$ is the inflation of $\mathcal{PB}(A,\emptyset)$ in the sense of Definition˜3.5. Precisely, for a vertex $\{[H]\}$ in $\mathcal{PB}(A,\emptyset)$, we take $P_{[H]}$ as the set of $A$-conjugacy classes of generators of $H$. Since $\mathcal{PB}(A,\emptyset)$ is Cohen–Macaulay of dimension n1n-1 by Theorem˜1.3, $\mathcal{B}(A)$ is also Cohen–Macaulay of dimension n1n-1 by [20, Proposition 5.10]. ∎